# Fairness Risks for Group-Conditionally Missing Demographics

### v2

@@ 3 Preliminary
Here, for a finite set $S$, we denote the mean of its elements as $\text{mean}\,(S)$. The above estimator is straightforward from Eq 1 and 7 of Menon and Williamson (2018) in the setting of randomized classifier. It is asymptotically unbiased, consistent, and differentiable in $f$; see the proof in Appendix E. Similarly, we can estimate $P(\hat{Y}=1|A=a,Y=y)\approx\text{mean}\,(S),{\quad\text{where}\quad}S:=\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=a,y_{i}=y\}.$ Other estimators can also be adopted, such as kernel density estimation (Cho et al., 2020). Applying these estimates to MD, DI, $\chi^{2}$-divergence, or any other metric, we obtain a fairness risk $\mathcal{F}(P_{f},\mathbf{\bm{y}},\mathbf{\bm{a}})$. We keep its expression general, and some concrete examples are given in (18) to (20) in Appendix A.1. An example based on (5) is
@@ 4.1 Rationalizing semi-supervised fairness risk
However, sharing backbones will allow $q_{\phi}(A|X,\tilde{A})$ to be influenced by the learning of $P_{f}(Y|X,\tilde{Y})$, conflicting with the aforementioned recusal principle. In light of this difficulty, we resort to the commonly used stop-gradient technique that is available in PyTorch and TensorFlow. While they still share backbones, the derivative of $\mathcal{E}_{\text{vanilla}}$ in (7) with respect to $q_{\phi}(A|X,\tilde{A})$ is stopped (set to 0) from backpropagation, i.e., treating $q_{\phi}(A|X,\tilde{A})$ as a constant. It also much simplifies the training process.
@@ 5.2 Instilling Fairness to SS-VAE
The evidence lower bound (ELBO) of $\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ can be derived in a standard fashion. Denoting it as $\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$, we can extend it as
$\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\geq\!\!\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{z}},a,y)\sim q_{\phi}(\cdot|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})}[\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}|a,y,\mathbf{\bm{z}})\!$
$+\log(p(y)p(a)p(\mathbf{\bm{z}}))-\log q_{\phi}(a,y,\mathbf{\bm{z}}|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})]$
$=\mathop{\mathbb{E}}\limits_{\mathbf{\bm{z}}\sim q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})}\log p_{\theta}(x|a,y,\mathbf{\bm{z}})+\mathop{\mathbb{E}}\limits_{y\sim q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})}\log P_{\theta}({\tilde{y}}|y)$
$+\mathop{\mathbb{E}}\limits_{a\sim q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})}\log P_{\theta}({\tilde{a}}|a)-\text{KL}[q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})||p(\mathbf{\bm{z}})]$
$-\text{KL}[q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})||p(y)]-\text{KL}[q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})||p(a)]$
$=:\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}).$
Followed by ELBO, we extend the SS-VAE objective as follows, to be minimized over $\theta$ and $\phi$:
@@
$-\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\sim\mathcal{D}_{tr}}\!\!\Big{[}\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ (12)
$+[\![{\tilde{a}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{a}\text{Cat}(a|h_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{a}}|a)$
$+[\![{\tilde{y}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{y}\!\!\text{Cat}(y|g_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{y}}|y)\Big{]}.$
Here, the regularizer $\Omega$ enforces the Dirichlet prior. Suppose ${\tilde{y}}\neq\emptyset$. If the provided ${\tilde{y}}$ always truthfully represents $y$, then $P({\tilde{y}}|y)=[\![{\tilde{y}}=y]\!]$. As a result, the last line in (12) equals $\text{Cat}({\tilde{y}}|g_{\phi}(\mathbf{\bm{x}}))$, the standard supervised loss in SS-VAE.
@@
where $\lambda>0$ is a tradeoff hyperparameter. We will henceforth refer to this model as Fair-SS-VAE.
@@ 6 Experimental Results
CelebA (Liu et al., 2015). CelebA contains face images annotated with 40 binary attributes. We sampled 45k images as our training and validation dataset, and 5k as test set. Following Jung et al. (2022), we set ”Attractive” as the target label, and ”Gender” as the sensitive group attribute.
@@
a
@@ 6.1 Main results
To investigate the significance of stop gradient, we set up an ablation experiment that directly uses $\mathcal{E}_{\text{vanilla}}$. The corresponding results are labeled “Ablation” in Figure 2, 3 and A.2. Under most circumstances, it performs worse than Fair-SS-VAE, with occasional advantages when applied to datasets with dense sparsity levels.
@@ A.3 Straight-through Gumbel-Softmax for vanilla fairness risk
A.4 Soften the step function to enable differentiation
@@
A.5 Labeling test data by minimizing fairness and classification risks
@@
A.6 Extension to class-probability estimation
@@ Appendix C Additional Experimental Results
Figure A.2 shows the Pareto frontier of error versus DEO/DEOPP for Adult-Race.

### v3

@@ 3 Preliminary
Here, for a finite set $S$, we denote the mean of its elements as $\text{mean}\,(S)$. The above estimator is straightforward from Eq 1 and 7 of Menon and Williamson (2018) in the setting of randomized classifier. It is asymptotically unbiased, consistent, and differentiable in $f$. see the proof in Appendix E. Similarly, we can estimate $P(\hat{Y}=1|A=a,Y=y)\approx\text{mean}\,(S),{\quad\text{where}\quad}S:=\{P_{f}(\hat{Y}=1|x_{i}):a_{i}=a,y_{i}=y\}.$ Other estimators can also be adopted, such as kernel density estimation (Cho et al., 2020). Applying these estimates to MD, DI, $\chi^{2}$-divergence, or any other metric, we obtain a fairness risk $\mathcal{F}(P_{f},\mathbf{\bm{y}},\mathbf{\bm{a}})$. We keep its expression general, and some concrete examples are given in (18) to (20) in Appendix A.1. An example based on (5) is
@@ 4.1 Rationalizing semi-supervised fairness risk
However, sharing backbones will allow $q_{\phi}(A|X,\tilde{A})$ to be influenced by the learning of $P_{f}(Y|X,\tilde{Y})$, conflicting with the aforementioned recusal principle. In light of this difficulty, we resort to the commonly used stop-gradient technique that is available in PyTorch (detach) and TensorFlow (tf.stop_gradient). While they still share backbones, the derivative of $\mathcal{E}_{\text{vanilla}}$ in (7) with respect to $q_{\phi}(A|X,\tilde{A})$ is stopped (set to 0) from backpropagation, i.e., treating $q_{\phi}(A|X,\tilde{A})$ as a constant. It also much simplifies the training process. The stop-gradient method is a technique that blocks the flow of gradient through a specific part of the computational graph during backpropagation, which is commonly used to decouple certain computations from parameter updates. We present the corresponding results in our ablation study, shown by the blue curve among all experiment figures.
@@ 4.3 Efficient evaluation of fairness risk $\mathcal{E}$
This theorem is non-trivial because the fairness risk can depend on the predicted probabilities in a complex manner, creating nonlinear dependencies among all the training examples; see Appendix A.1. The analysis also reveals the limitations of the sampler, as certain risks, such as disparate impact, are unbounded.
@@ 5 Integrating Fairness Risk with SSL
Why VAE?
In terms of SSL, while some discriminative methods exist such as Zhu et al. (2003), the most prevalent and effective approaches are based on generative models, such as SS-VAE. Indeed, most conventional generative models for SSL are essentially VAEs, especially when they are trained with variational inference.
Secondly, our approach to fairness risk relies on an encoder capable of estimating the posterior probability of missing labels or demographics. This rules out normalizing flows or generative adversarial networks unsuitable, as the latter’s discriminator does not serve as an encoder. Moreover, transformers are not considered since our formulation does not involve sequential data modeling.
In contrast, VAE employs an encoder and provides the conditional independencies that allow us to more finely specify the constituent distributions as in (10) and (5.1). Diffusion models, as a more refined variant of VAE, can also be adopted. However, we intentionally keep the generative model simple, prioritizing fairness modeling, while future work can incorporate additional refinements to the VAE.
@@ 5.2 Instilling Fairness to SS-VAE
The evidence lower bound (ELBO) of $\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ can be derived in a standard fashion, and we relegate the details to Appendix A.4. Denoting it as $\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$, we extend the SS-VAE objective as follows, to be minimized over $\theta$ and $\phi$:
@@
$\qquad\qquad-\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})\sim\mathcal{D}_{tr}}\!\!\Big{[}\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ (12)
$\qquad\qquad\qquad+[\![{\tilde{a}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{a}\text{Cat}(a|h_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{a}}|a)$
$\qquad\qquad\qquad+[\![{\tilde{y}}\neq\emptyset]\!]\cdot\log\sum\nolimits_{y}\!\!\text{Cat}(y|g_{\phi}(\mathbf{\bm{x}}))P_{\theta}({\tilde{y}}|y)\Big{]}.$
Here, the regularizer $\Omega$ enforces the Dirichlet prior. For example, let $a=P(\text{male}|\text{male})$, $b=P(\text{female}|\text{male})$, $c=P(\emptyset|\text{male})$. Then a prior of $\text{Dir}(0.5,0.2,0.4)$ leads to a regularizer of $\log(a^{0.5-1}b^{0.2-1}c^{0.4-1})$. The regularization is beneficial when prior knowledge about the distributions is available, such as the likelihood that a male is more likely to be labeled as male rather than female. However our experiments did not make any such assumptions and thus did not incorporate this regularizer. Suppose ${\tilde{y}}\neq\emptyset$. If the provided ${\tilde{y}}$ always truthfully represents $y$, then $P({\tilde{y}}|y)=[\![{\tilde{y}}=y]\!]$. As a result, the last line in (12) equals $\text{Cat}({\tilde{y}}|g_{\phi}(\mathbf{\bm{x}}))$, the standard supervised loss in SS-VAE.
@@
where $\lambda>0$ is a tradeoff hyperparameter and $\mathcal{E}$ takes the same form as in (8). We will henceforth refer to this model as Fair-SS-VAE.
@@ 6 Experimental Results
CelebA (Liu et al., 2015). CelebA contains face images annotated with 40 binary attributes. We sampled 45k images as our training and validation dataset, and 5k as test set. Following (Jung et al., 2022), we set ”Attractive” as the target label, and ”Gender” as the sensitive group attribute.
@@
It is worth mentioning that conducting experiments where there is a significant correlation between target labels and groups is necessary. If the labels and groups are already uncorrelated, fairness modeling is likely less critical. For example, in CelebA, when considering a label/group pair with very low correlation, such as ’Eyeglasses’ and ’Male’, the model already achieves high accuracy with a DEO close to zero (ranging from 0.000 to 0.003) even without any fairness interventions. This suggests that fairness-aware models, including ours, have limited potential for improvement, making them less ideal for experiments.
@@
Precisely, we focused on varying the tradeoff weight between accuracy and fairness while keeping other hyperparameters fixed, such as batch size, learning rate, and the weights for KL divergence and reconstruction error in the VAE. Since these hyperparameters are already part of the standard SS-VAE, we relied on SS-VAE to select their values. After determining these values, we put back the fairness risk and further adjusted the selected hyperparameters, generating a new set of candidate values for model selection based on the validation set.
@@ 6.1 Main results
To investigate the significance of stop gradient, we set up an ablation experiment that directly uses $\mathcal{E}_{\text{vanilla}}$. The corresponding results are labeled “Ablation” in Figure 2, 3 and A.2. Under most circumstances, it performs worse than Fair-SS-VAE, with occasional advantages when applied to datasets with dense sparsity levels. Additionally, we carried out experiments titled CelebA-HighCheekbones, as illustrated in E, where high cheekbones is the target label and gender serves as the group attribute.
@@
Acknowledgement
We thank the reviewers and the meta-reviewer for assessing our paper and for their constructive feedback. This work is supported by NSF grant RI:1939743.
@@ A.3 Straight-through Gumbel-Softmax for vanilla fairness risk
A.4 ELBO
We can extend the ELBO as
$\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})$ $\geq\!\!\mathop{\mathbb{E}}\limits_{(\mathbf{\bm{z}},a,y)\sim q_{\phi}(\cdot|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})}[\log p_{\theta}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}|a,y,\mathbf{\bm{z}})\!+\log(p(y)p(a)p(\mathbf{\bm{z}}))-\log q_{\phi}(a,y,\mathbf{\bm{z}}|\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}})]$
$=\mathop{\mathbb{E}}\limits_{\mathbf{\bm{z}}\sim q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})}\log p_{\theta}(x|a,y,\mathbf{\bm{z}})+\mathop{\mathbb{E}}\limits_{y\sim q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})}\log P_{\theta}({\tilde{y}}|y)+\mathop{\mathbb{E}}\limits_{a\sim q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})}\log P_{\theta}({\tilde{a}}|a)$
$-\text{KL}[q_{\phi}(\mathbf{\bm{z}}|\mathbf{\bm{x}})||p(\mathbf{\bm{z}})]-\text{KL}[q_{\phi}(y|\mathbf{\bm{x}},{\tilde{y}})||p(y)]-\text{KL}[q_{\phi}(a|\mathbf{\bm{x}},{\tilde{a}})||p(a)]$
$=:\text{ELBO}(\mathbf{\bm{x}},{\tilde{a}},{\tilde{y}}).$
A.5 Soften the step function to enable differentiation
@@
A.6 Labeling test data by minimizing fairness and classification risks
@@
A.7 Extension to class-probability estimation
@@ Appendix C Additional Experimental Results
Figure A.2 shows the Pareto frontier of error versus DEO/DEOPP for Adult-Race. Figure E shows the results for Celeba-HighCheekbones. Here, we use high cheekbones as the label and gender as the group attribute.
@@ Appendix E Statistical Property of Estimator (5)
(a) DEO-Sparse
(b) DEO-Medium
(c) DEO-Dense
(d) DEOPP-Sparse
(e) DEOPP-Medium
(f) DEOPP-Dense
Figure 8: Pareto frontier of error versus DEO/DEOPP for CelebA-HighCheekbones